# references.bib

### v1

@@ I Introduction
However, the SDE in endoscopic imagery remains largely domain-specific, suffering from poor generalization and limited real-world application. Despite some efforts to adapt foundation models trained on natural imagery for medical depth estimation, these are limited in scope and primarily experimental [cui2024surgical, cui2024endodac, han2024depth]. The disparity in research focus between natural and medical imaging can be attributed to two challenges inherent in medical datasets: 1) data scarcity and 2) inferior labeling quality [gonzalez2023robust]. These factors result in models that are prone to overfitting and struggle to generalize to unseen data [zhon2024robust].
@@
CholecT50 [azagra2023endomapper] Abdomen 100,863 fr.
@@ III-A Teacher Model Training
Several endoscopic datasets with depth annotations [allan2021stereo, bobrow2023, edwards2020serv, ozyoruk2021endoslam, penza2018endoabs, rau2019implicit] are employed. Accurately annotating depth labels in endoscopic data presents a significant challenge compared to general open scenes because high-precision sensors are impractical to set up during clinical procedures. Consequently, existing datasets generate annotations using stereo cameras [allan2021stereo, edwards2020serv, penza2018endoabs], registration to CT scans [bobrow2023, ozyoruk2021endoslam], or synthetic data [ozyoruk2021endoslam, rau2019implicit]. To expand the training data, we have curated an additional dataset for bronchoscopy. Depth annotations are sourced using registration-based methods, airway meshes from patient CT scans are reconstructed, and video frames are manually registered to these in a virtual endoscopic view. Depth labels are generated by rendering depth maps in the airway model with the registered endoscopic poses. As a result, our teacher model leverages these extensive datasets, further enhancing training data diversity by generating pseudo-labels for unlabeled datasets to guide the student model training.
@@ III-B Robust Student Model Training
To extend training data diversity for student model, besides labeled data from sources [allan2021stereo, bobrow2023, edwards2020serv, ozyoruk2021endoslam, penza2018endoabs, rau2019implicit], we have incorporated unlabeled endoscopic data from datasets intended for other tasks [azagra2023endomapper, nwoye2022rendezvous, bernal2015wm, zia2021surgical, ross2021comparative, ali2020objective]. Each dataset introduces unique challenges due to variations in annotation accuracy—depth labels from stereo cameras can be sparse and imprecise due to calibration errors and environmental factors, while labels from registration methods may suffer from errors in registration and CT reconstruction inaccuracies. Leveraging this mix of unlabeled data enhances the breadth of our training data, yet introduces the risk of noisy pseudo-labels from the teacher model, which can mislead the student model’s training. This is particularly problematic in medical imaging, where fewer labeled data are available to train the teacher model, leading to its unstable performance on out-of-distribution unlabeled data. Such noise in training could result in degraded generalization performance during practical application. To address this, we propose to mitigate the impact of these inaccuracies by estimating the confidence of each labeled pixel. We then adjust the learning weight for each pixel based on its estimated confidence. This approach focuses gradient updates on regions with higher confidence, while reducing the influence of gradients from noisier regions.
@@
III-C Removing the Ambiguous Shift Factor
To ensure a fair comparison with the latest research in medical image depth estimation, we fine-tune our foundation model $f$ to create a refined model $\dot{f}$ for zero-shot relative depth estimation (RDE), where only an unknown scale factor remains. Note that our main contribution remains the foundation model $f$. To address the shift ambiguity, we introduce a simple but effective third step: fine-tuning the student model $f$ with a frozen encoder on the SCARED dataset using scale-invariant loss supervision. The SCARED dataset was chosen for its high annotation accuracy and sufficient training frames for fine-tuning. Formally, the pretrained student model $f$ is fine-tuned on a single dataset using scale-invariant loss [nips2014_7bccfde7]:
@@ IV-A Implementation Details
Cycle-GAN [zhu2017unpaired] 0.433 3.460 7.091 0.407 0.355
DD-VNB [tian2024dd] 0.715 11.997 9.916 0.528 0.367
@@
  • $\ddagger$ denotes in-domain evaluation.
@@ IV-D Fine-tuned Metric Depth Estimation
Zero-Shot Metric Depth Estimation. We further evaluate the zero-shot MDE performance of our model on phantom bronchoscopy data. We choose the model fine-tuned on our patient bronchoscopy data for their similar depth range, making it possible for zero-shot generalization. The evaluation test set of phantom data aligns with data used in [tian2024dd]. As shown in Table VI, our model achieves better zero-shot MDE performance even when compared to methods that are tested on in-domain data. Furthermore, we compare using Depth Anything [yang2024depth] pretrained weights to initialize the MDE model encoder and perform zero-shot MDE on phantom bronchoscopy data in Table VI. By training with our robust loss on abundant mixed endoscopic data, our model learns better endoscopy image representations and achieves better zero-shot MDE performance on all evaluated metrics.

### v2

@@ I Introduction
However, the SDE in endoscopic imagery remains largely domain-specific, suffering from poor generalization and limited real-world application. Despite some efforts to adapt foundation models trained on natural imagery for medical depth estimation, these are limited in scope and primarily experimental [cui2024surgical, cui2024endodac, han2024depth]. The disparity in research focus between natural and medical imaging can be attributed to two challenges inherent in medical datasets: 1) data scarcity [cui2020unified] and 2) inferior labeling quality [xue2022robust]. These factors result in models that are prone to overfitting and struggle to generalize to unseen data [zhon2024robust].
@@
CholecT50 [nwoye2022rendezvous] Abdomen 100,863 fr.
@@ III-A Teacher Model Training
Several endoscopic datasets with depth annotations [allan2021stereo, bobrow2023, edwards2020serv, azagra2023endomapper, penza2018endoabs, rau2019implicit] are employed. Accurately annotating depth labels in endoscopic data presents a significant challenge compared to general open scenes because high-precision sensors are impractical to set up during clinical procedures. Consequently, existing datasets generate annotations using stereo cameras [allan2021stereo, edwards2020serv, penza2018endoabs], registration to CT scans [bobrow2023], or synthetic data [azagra2023endomapper, rau2019implicit]. To expand the training data, we have curated an additional dataset for bronchoscopy. Depth annotations are sourced using registration-based methods, airway meshes from patient CT scans are reconstructed, and video frames are manually registered to these in a virtual endoscopic view. Depth labels are generated by rendering depth maps in the airway model with the registered endoscopic poses. As a result, our teacher model leverages these extensive datasets, further enhancing training data diversity by generating pseudo-labels for unlabeled datasets to guide the student model training.
@@ III-B Robust Student Model Training
To extend training data diversity for student model, besides labeled data from sources [allan2021stereo, bobrow2023, edwards2020serv, penza2018endoabs, rau2019implicit, azagra2023endomapper], we have incorporated unlabeled endoscopic data from datasets intended for other tasks [ozyoruk2021endoslam, azagra2023endomapper, nwoye2022rendezvous, bernal2015wm, zia2021surgical, ross2021comparative, ali2020objective]. Each dataset introduces unique challenges due to variations in annotation accuracy—depth labels from stereo cameras can be sparse and imprecise due to calibration errors and environmental factors, while labels from registration methods may suffer from errors in registration and CT reconstruction inaccuracies. Leveraging this mix of unlabeled data enhances the breadth of our training data, yet introduces the risk of noisy pseudo-labels from the teacher model, which can mislead the student model’s training. This is particularly problematic in medical imaging, where fewer labeled data are available to train the teacher model, leading to its unstable performance on out-of-distribution unlabeled data. Such noise in training could result in degraded generalization performance during practical application. To address this, we propose to mitigate the impact of these inaccuracies by estimating the confidence of each labeled pixel. We then adjust the learning weight for each pixel based on its estimated confidence. This approach focuses gradient updates on regions with higher confidence, while reducing the influence of gradients from noisier regions.
@@ IV-A Implementation Details
For a fair comparison with recent work in medical image depth estimation, we fine-tune our foundation model $f$ to produce a refined model $\dot{f}$ for zero-shot relative depth estimation (RDE), leaving only an unknown scale factor. Note that the key contribution remains the foundation model $f$. We fine-tune the student model $f$ with a frozen encoder on the SCARED dataset using scale-invariant loss. SCARED was selected for its high annotation accuracy and sufficient training frames. Formally, the scale-invariant loss[nips2014_7bccfde7] is defined as:
@@
Cycle-GAN [zhu2017unpaired] 0.583 6.427 9.953 0.488 0.234
DD-VNB [tian2024dd] 0.498 4.064 7.628 0.424 0.331
@@
  • All methods trained on patient data and tested on phantom data.
@@ IV-D Fine-tuned Metric Depth Estimation
Zero-Shot Metric Depth Estimation. We further evaluate the zero-shot MDE performance of our model on phantom bronchoscopy data. We choose the model fine-tuned on our patient bronchoscopy data for their similar depth range, making it possible for zero-shot generalization. The evaluation test set of phantom data aligns with data used in [tian2024dd]. As shown in Table VI, our model achieves better zero-shot MDE performance compared to SOTA methods on all evaluated metrics with a large margin. Furthermore, we compare using Depth Anything [yang2024depth] pretrained weights to initialize the MDE model encoder and perform zero-shot MDE on phantom bronchoscopy data in Table VI. By training with our robust loss on abundant mixed endoscopic data, our model learns better endoscopy image representations and achieves better zero-shot MDE performance on all evaluated metrics.